Bill can row 3mph in still water. It takes him 3 hours 36 minutes to go 3 miles upstream and return. Find the speed of the current.
Answer by lwsshak3(11628)

Bill can row 3mph in still water. It takes him 3 hours 36 minutes to go 3 miles upstream and return. Find the speed of the current.
change 3 hours and 36 minutes to 3 and 3/5 of an hour = 3.6 hour
let x=speed of current
downstream speed = 3+x
upstream speed = 3-x
time=distance/speed
3/(3+x)+3/(3-x)=3.6
3(3-x)+3(3+x) =3.6(3-x)(3+x)
9-3x+9+3x=3.6(9-x^2)
18=32.4-3.6x^2
3.6x^2=32.4-18=14.4
x^2=14.4/3.6=4
x=sqrt(4)=2 mph or-2 mph(reject)
ans: The speed of the current is 2 mph